OpenStudy (anonymous):

Which value of a would make the following statement true? 6(2x + a) = 12x + 30 5 -5 4 -4

OpenStudy (anonymous):

Again, you need to use the distributive property 12x + 6a = 12x + 30

OpenStudy (anonymous):

6a = 30 a = 5

OpenStudy (anonymous):

Thanks @KirbyLegs :)

OpenStudy (shamil98):

You don't need to use the distributive property.. simply take the gcf of the right side and divide :) 6(2x + a) = 12x + 30 6(2x + a) = 6(2x+5) divide by 6 2x + a = 2x +5 subtract 2x from both sides. a = 5

OpenStudy (anonymous):

I didn't think about it that way. Thanks @shamil98